OpenStudy (anonymous):

Find each function value, if it exists for f(t) = sqrt t^2+1. f(-1)?

OpenStudy (across):

Tell me, what is confusing you about this problem?

OpenStudy (anonymous):

Well I know there is no square root of a negative number, right?

OpenStudy (anonymous):

So I came up with zero as my answer but it says that is not correct.

OpenStudy (across):

I will assume the function is the following:\[f(t)=\sqrt{t^2+1}.\]What happens if you pass it a -1? I.e.,\[f(-1)=\sqrt{(-1)^2+1}?\]

OpenStudy (across):

By the way, you're right that you can't take the square root of a negative number (that is, if you're not working with complex numbers).

OpenStudy (anonymous):

My other option is that the square root is not a real number and that isn't right either, so I am stumped.

OpenStudy (across):

\[f(-1)=\sqrt{(-1)^2+1}=\sqrt{1+1}=\sqrt{2}.\] Do you agree?

OpenStudy (anonymous):

How did you get (-1)^2 = 1? I figured it as being -2

OpenStudy (across):

Because\[(-1)^2=(-1)(-1)=1.\]A negative number times a negative number is always a positive number.

OpenStudy (anonymous):

okay, well that is where I am messing up.

OpenStudy (anonymous):

I also forgot to put the "sqrt" symbol in front of my answer. That was stupid!

OpenStudy (across):

No, that's fine; we all forget a thing or two at times. ;]

OpenStudy (anonymous):

Thanks for your help!